OpenStudy (anonymous):

Find the solution: -3 sin^-1 (2x) = π

hartnn (hartnn):

can you isolate sin^-1 (2x) first ?

OpenStudy (anonymous):

It equals -π/3?

hartnn (hartnn):

yes, now if \(\sin^{-1}a=b, \implies a=\sin b\) so, what can you say baout sin^-1 (2x) =-pi/3 ?

OpenStudy (anonymous):

sin(π/3)=2x?

hartnn (hartnn):

what about -*minus* sign ? why did you just drop it ? :P

hartnn (hartnn):

won't it be sin(-π/3)=2x

OpenStudy (anonymous):

Oops, I forgot it, sorry! sin(-π/3)=2x

hartnn (hartnn):

you can now use \(\sin (-\theta )=-\sin \theta\) to get -sin(π/3)=2x got this ? and whats sin (pi/3)=...?

OpenStudy (anonymous):

I'm not sure, would it be √3)/2?

hartnn (hartnn):

yes, thats correct. so, -(√3)/2=2x i believe you can find 'x' from here easily :)

OpenStudy (anonymous):

You get -((√3)/4)?

hartnn (hartnn):

thats correct value of x :)

OpenStudy (anonymous):

Thank you again! (:

hartnn (hartnn):

Welcome again! (: